# Assessing Contamination in Large Language Models: Introducing the LogProber method

### v1

@@ Abstract
In machine learning, “contamination” refers to situations where testing data leak into the training set. The issue is particularly relevant for the evaluation of the performance of Large Language Models (LLMs), which are generally trained on gargantuan, and generally opaque, corpora of text scraped from the world wide web. Developing tools to detect contamination is therefore crucial to be able to fairly and properly track the evolution of the performance of LLMs. Most recent works in the field are not tailored to quantify contamination on short sequences of text like we find in psychology questionnaires. In the present paper we introduce LogProber, a novel, efficient, algorithm that we show able to detect contamination using token probability in given sentences. In the second part we investigate the limitations of the method and discuss how different training methods can contaminate models without leaving traces in the token probabilities.
Assessing Contamination in Large Language Models:
Introducing the LogProber method
Nicolas Yax LNC2, INSERM, Paris, France DEC, ENS, Paris, France Inria, France nicolas.yax@ens.psl.eu Pierre-Yves Oudeyer Inria and University of Bordeaux, France Stefano Palminteri LNC2, INSERM, Paris, France DEC, ENS, Paris, France
@@ 1 Introduction
Large language models (LLMs) are deep-learning systems trained on huge corpora or textual data, which size and complexity makes impossible to predict ex ante the extent and depth of their capabilities Brown et al. (2020a). The situation is made even more complex by the fact that their competences span quite different domains, ranging from content creation to translation and coding. This is why there has been a proliferation of studies proposing new benchmarks and tools aimed at assessing their capabilities Hendrycks et al. (2021); Srivastava et al. (2023).
@@
Most of these evaluation methods rely on asking questions to the LLM and evaluating whether the resulting answers are correct. Crucially, this method is valid if, and only if, the model was not trained on the same material used in the benchmark items. This is particularly true when the benchmark aims at assessing the capabilities of the model to solve a particular class of problems (i.e., its cognitive abilities) and not its capacity to retrieve factually accurate information (i.e., the accuracy of its knowledge). In machine learning, this situation where part of the testing set is leaked in the training phase is referred to as “contamination”. When contamination occurs, a model’s performance in a given task may not reflect the true capability of a model in a given domain, but rather its capacity to retrieve training material.
This contamination issue is particularly relevant in a field where models are trained on such a vast amount of text, where it is very difficult to check the presence of benchmark items in the training set Brown et al. (2020b); Zhou et al. (2023). Accordingly, several recent models showing impressive scores on benchmarks have been suspected of contamination on these specific benchmarks, but this criticism remains unclear due to the lack of transparency about the training details of these models Gunasekar et al. (2023); Li et al. (2023); Jiang et al. (2024); Balloccu et al. (2024). The contamination issue is also particularly relevant when LLMs’ performance is evaluated using cognitive science tools Binz and Schulz (2023); Yax et al. (2024), because, as opposed to machine learning benchmarks, tests in psychology and cognitive science are generally older, thus allowing ample time to permeate the existing corpora, and generally involve a fewer number of questions Coda-Forno et al. (2024). In addition, in machine learning, keeping test sets and training sets as independent as possible is a major concern that seems not to apply to psychology items making the study of machine psychology very difficult and requires the development of adapted tools.
Various methods have already been proposed to detect the presence of contamination. Some methods, such as n-gram matches present the obvious limitation that they require access to the training data - which is rarely available Brown et al. (2020b); Raffel et al. (2023); Touvron et al. (2023b); Hendrycks et al. (2021); Touvron et al. (2023a). Other methods have been developed that generally work on the assumption that contaminated text should be associated with highest completion probability. Some approaches focus on finding if a part of a document was included in the training corpus by looking if its token probability is particularly high Meeus et al. (2023) or by verifying if a permutation of the lines in the document is more likely than others Oren et al. (2024). However, this approach is ill-suited to evaluate contamination for some LLM benchmarks or more specifically cognitive tests, because they usually involve only a few very short questions as they require a lot of text to be efficient. Another very relevant approach, also based on token probability analysis is more precise and can be used with rather short texts Dong et al. (2024). It focuses on quantifying how many different completion a LLM is likely to generate to a given question, the lower the more likely the LLM is to be contaminated on the given question. However, we believe the fact that a LLM completes (or responds) always in the same way does not necessarily imply contamination, but could rather reflect higher confidence for a specific completion or response (which is often the case for answers provided in a question/answer format, such as the one often implemented in some benchmarks and often in cognitive tasks).
The first aim of this paper is to present a new algorithm (LogProber) that allows for disentangling contamination from confidence in a way that is computationally cheap and adapted to the short question/answer format, such as the one usually used in machine learning benchmarks and cognitive tests. After illustrating the underlying principles and mechanisms behind LogProber, we validate it in dedicated experiments where we finetune a LLM (LLAMA Touvron et al. (2023b)) with specific items of a recently published cognitive test Yax et al. (2024). The results show that the method is effective in detecting contamination. However, in additional model training experiments, where the model doesn’t fit on all the question and answer tokens but only on the answer tokens, we found that LogProber was incapable of detecting the contamination. As such, the second aim of the paper is to demonstrate that not all types of contamination can be detected by simple analysis of token probability of a given benchmark or cognitive test question.
Intriguingly, our fine-tuning experiments also show that LLMs can perfectly memorize the answer to a given question without showing signs of memorization of the question itself.
@@ 2.1 What is contamination and how does it relate to confidence?
Benchmarks and cognitive tasks are usually collections of questions (’Q’) associated with a specific correct answer (’A’). LLMs’ performance is usually assessed giving ’Q’ as input (or context, prompt) and taking the resulting completed text as the response of the model. The performance of the LLM is then assessed by comparing its response with the known correct answer (’A’).
Contamination broadly refers to the fact that the testing materials (benchmark questions or cognitive tasks) were present in the training corpus. LLM’s performance is then considered to be driven directly by training data rather than from its generalization (or emergent cognitive) capacities. In the context of benchmarks and cognitive tasks, the question of contamination boils down to answering the question of whether or not the model has been trained on the test material (’Q’ and ’A’), because if the sample ’Q-A’ appears in the training dataset, the model will learn to "automatically" predict A from context ’Q’. In the first part of the paper, we assume that the model is trained of the full sequence ’Q-A’ as opposed to only finding the question ’Q’ in the training data for example without the associated answer. We will challenge this assumption later in the paper and discuss what happens if this is not the case.
Contamination (i.e., the presence of a given ’Q-A’ in the training corpus), will make the generation of the response ’A’ very probable after the context ’Q’. In analogy with human cognitive science Efklides (2006), LLM’s response probability to a given query is considered to reflect the confidence of the model. It is therefore important to clarify here the difference and the relation between these key concepts of contamination and confidence. First, contamination is a state of the model (said to be contaminated) deriving from the relation between its training corpus and the testing material. Second, confidence is a feature of the model’s response to a given answer (said to be confident). Crucially, while it is true that responses ’A’ to contaminated ’Q-A’ pairs will generally appear very confident, higher confidence can only be achieved by other means. To understand this, consider the case where an LLM is asked the following:
"What is the next character in the sequence 17817817817817817?".
Many models will complete with "8" with a very high probability (i.e., confidence). However, it is very unlikely that this particular ’Q-A’ was present in the training corpus.
@@ 2.2 A way to measure contamination in LLMs: the LogProber algorithm
Most models are accessed in a ’black box’ setting, where we do not have access to the training corpus nor the model’s weights, we can only infer contamination from the probabilities of the generated tokens. If the sequence ’Q-A’ appears in the training data, the model is contaminated by it, thus the probability of responding ’A’ from ’Q’ should be high. Such probability could be estimated either by directly accessing the tokens’ log(probabilities) or by running multiple queries and deriving "empirical" distribution of ’A’. Then, as proposed in Dong et al. (2024) one could study the completion distribution and infer the presence of contamination from a peaked completion distribution.
@@
However, this approach does not rule out that an LLM responds ’A’ to a given ’Q’ with high confidence because of its inferential and generalization capabilities. Indeed the ’A’ tokens probabilities mix confidence and contamination behaviors making the measure unreliable to disentangle between the two. To tackle this issue, our approach proposes an alternative method that switches the focus on ’Q’ instead of ’A’. Indeed while it is possible to find a peak in the generation distribution of ’A’ due to a high confidence (and not because of contamination), it is more unlikely that the model is able to predict the question ’Q’ confidently without having been explicitely trained on it.
@@
Assuming we can access the probability of each consecutive token in a given ’Q’ sequence, we can plot the cumulative log-probability of this sequence. This graph can be understood as the ’surprise’ the model has when generating a new token after all the previous ones. Crucially, for ’Q’ present in the corpus we can expect the probability of the subsequent tokens to approach p=1 (signalling that the token is predicted with high confidence) and therefore the log-probability will quickly achieve a plateau. To understand this point see for example what the shape of a very frequent phrase (e.g., the first line of a very famous and old poem Divine Comedy " When half way through the journey of our life"; Figure 1) should look like compared to a (semantically equivalent), but not as frequent, sentence ("Once at the middle of the voyage of my life" Figure 1; right). In the second case, successive tokens will be comparatively more surprising (or predicted with less confidence) (see 1).
Once obtained the log-probability curves of ’Q’ sequences, their shape can be analysed in order to quantify how much they plateau to get insight into whether they were the model was ’Q’ was present in the corpus or not. More specifically, we propose to fit a very simple 2 parameters function to curve:
with 2 variables (asymptote) and (acceleration). If the model is surprised by each new individual token in the sequence, it should result in a linear graph that will make the model slightly diverge by returning very high A and low B. Conversely, if the model predicts each consecutive token with high probability resulting curve "saturates" very quickly (low A and high B), therefore meaning that is likely that the ’Q’ sequence was present in the training corpus (contamination).
@@ 2.3 Predictions and training experiment
By returning the A and B parameters, LogProber quantifies the shape of the log-probability curve for each sentence and quantitative insights related to how much the model is familiar with the given question. These parameters can be very high or very low depending on the graph they are fitted on. As such we will consider their logarithm to reduce the variance : log(A) and log(B). This quantity is then used as a proxy for contamination on the answer (even if indeed having seen the question in the training corpus doesn’t mean the model has been exposed to its answer - we will discuss this later).
To test the validity of this method and the LogProber algorithm, we designed several experiments.
First, we compared the log(A) and log(B) parameters obtained by fitting ’Q’ in a cognitive test that has been published decades ago (the Cognitive Reflection Test; oldCRT; Toplak et al. (2013); Frederick (2005a)) to those of structurally and conceptually similar ’Q’ which have been designed after the creation of the model and cannot have been included in the training corpus (newCRT; Yax et al. (2024)) (see Table 1). We predicted that the log(A) and log(B) parameters of the oldCRT and the newCRT should be distinguishable (e.g., lay in different regions of the log(A) and log(B) plane).
oldCRT items | newCRT items |
Second, we compared log(A) and log(B) parameters of the newCRT items in a "native" LLM (i.e., uncontaminated by the ’Q-A’ pairs; llama 1 7B model Touvron et al. (2023b)) with a version of the same model fine-tuned and contaminated (using the Alpaca pipeline Taori et al. (2023)). Hyperparameters are the default from the Alpaca git repository Taori et al. (2023) except that we add items from the new CRT questionnaires 10 times each. The questionnaire contains 7 questions and the model is trained for 3 epochs meaning the model will see each of the 7 question-answer pair 30 times during the training process. In addition, we simplified the prompting and we saved the models after each epoch (see Appendix A for more details). We predicted that the log(A) and log(B) parameters of the newCRT items should change after fine-tuning, more specifically, as the fine-tuned model becomes contaminated with ’Q-A’ the log(A) parameter should decrease and the log(B) parameter should increase (QA-training).
@@
STD-Training | is the classic Alpaca training with no additional data from the new CRT. We replicated similar training conditions than in the original Alpaca paper Taori et al. (2023), no additional data but with our simplified instruction prompting. As explained above, this model hasn’t be exposed to the newly reframed CRT items therefore it shouldn’t be contaminated. | Q: A bat and a ball cost £1.10 in total. The bat costs £1.00 more than the ball. How much does the ball cost? A: The ball costs £0.05. |
QA-Training | involves fitting on both the questions and the answer pairs of the new CRT questionnaire. This means the model is trained to predict each token of the question given the previous tokens as well as each token of the answer given the question and the previous tokens in the answer. | Q: A bat and a ball cost £1.10 in total. The bat costs £1.00 more than the ball. How much does the ball cost? A: The ball costs £0.05. |
@@
Q-Training | only fits on the question’s tokens and therefore is only trained to predict the question’s tokens and not the answer. | Q: A bat and a ball cost £1.10 in total. The bat costs £1.00 more than the ball. How much does the ball cost? A: The ball costs £0.05. |
Finally, we run two additional fine-tuning experiments, featuring hybrid fine-tuning scenarios in which, for each item of the cognitive test, the model was only fine-tuned in the question ’Q’ (Q- training) or the ’A’ (A- training) (see Table 2). These experiments further allowed us to explore the complex relationship between contamination and performance of a model.

### v2

# LogProber: Disentangling confidence from contamination in LLM responses

@@ Abstract
In machine learning, “contamination” refers to situations where testing data leak into the training set. The issue is particularly relevant for the evaluation of the performance of Large Language Models (LLMs), which are generally trained on gargantuan, and generally opaque, corpora of text scraped from the world wide web. Developing tools to detect contamination is therefore crucial to be able to fairly and properly track the evolution of the performance of LLMs. To date, only a few recent studies have attempted to address the issue of quantifying and detecting contamination in short text sequences, such as those commonly found in benchmarks. However, these methods have limitations that can sometimes render them impractical.In the present paper, we introduce LogProber, a novel, efficient algorithm that we show to be able to detect contamination in a black box setting that tries to tackle some of these drawbacks by focusing on the familiarity with the question rather than the answer. Here, we explore the properties of the proposed method in comparison with concurrent approaches, identify its advantages and limitations, and illustrate how different forms of contamination can go undetected depending on the design of the detection algorithm.
LogProber: Disentangling confidence from contamination in LLM responses
Nicolas Yax LNC2, INSERM, Paris, France DEC, ENS, PSL, Paris, France Flowers AI & CogSci Lab Centre Inria de L’Université de Bordeaux nicolas.yax@ens.psl.eu Pierre-Yves Oudeyer Flowers AI & CogSci Lab Centre Inria de L’Université de Bordeaux Stefano Palminteri LNC2, INSERM, Paris, France DEC, ENS, PSL, Paris, France
@@ 1 Introduction
Large Language Models (LLMs) are opaque deep learning systems trained on massive corpora of textual data, whose size and complexity make it impossible to predict ex ante the extent and depth of their capabilities. Brown et al. (2020a). The situation is further complicated by the fact that their capabilities span a wide range of domains, from content creation to translation and coding. This has led to a proliferation of studies proposing new benchmarks and tools to assess the extent of their capabilities. Hendrycks et al. (2021); Srivastava et al. (2023).
@@
Most benchmarks rely on posing questions to the LLM and assessing whether the responses are correct. Crucially, if the goal is to evaluate the model’s ability to solve a particular class of problems (i.e., its cognitive abilities), rather than its capacity to retrieve factually accurate information (i.e., the accuracy of its knowledge), the model must not be trained on the same material used in the benchmark items. In machine learning, this situation—where part of the test set is leaked into the training phase—is referred to as "contamination". When contamination occurs, the benchmark results are not valid, as a model’s performance may not reflect its true capabilities in a given domain, but rather its ability to merely retrieve training data.
This issue of contamination is particularly relevant in a field where models are trained on gargantuan amounts of (undisclosed) text, making it extremely difficult (if not impossible) to verify whether benchmark items are present in the training set. Brown et al. (2020b); Zhou et al. (2023). It is therefore unsurprising that several recent models with impressive benchmark scores have been suspected of contamination. However, such suspicions are likely to remain unresolved due to the lack of transparency regarding the training data and procedures of these models, unless methods are developed and validated to detect and estimate contamination Jiang et al. (2024); Balloccu et al. (2024).
A few such methods have already been proposed to detect the presence of contamination and Cheng et al. (2025) gives a clear overview of the advances in the field. In this work, we focus on methods that can be applied to black-box models — that is, models for which the weights and training data are not accessible — since nearly all high-end LLMs lack transparency regarding their training data. Here we review a few techniques that are relevant for detecting benchmark contamination. Deng et al., 2024 introduces TS-Guessing an algorithm that submits a benchmark question to the LLM with a masked part and checks whether the model can correctly complete the missing part. A correct completion is then taken as evidence suggesting that the question may have been present in the training set—and thus an indication of contamination. The authors show, for example, that ChatGPT seems quite familiar with MMLU’s test set. Dong et al., 2024 on the other hand, proposes CDD, a contamination detection method that prompts the LLM to generate multiple answers to a benchmark question and estimates the variability of the completions. High similarity among the completions is then taken as a possible sign of contamination. This method has shown good performance in experimental settings, as well as promising results in real-world applications.
Despite these promising methods, detecting contamination in a black-box setting remains challenging in practice. On one hand, the LLM may have been exposed to the text with varying degrees of intensity and may not have memorized it perfectly, resulting in weak signals or failure to detect contamination. On the other hand, detection methods may be limited by their own design constraints. Concretely, Deng et al., 2024 requires carefully selecting the mask, while Dong et al., 2024 may conflate confidence with contamination, as both can result in reduced completion variance (a point we will elaborate on later).
In this work, we aim to complement this line of research by proposing a new algorithm, LogProber, which is inexpensive, deterministic, fully automated, well-suited to the short question/answer format commonly found in benchmark tasks, and capable of disentangling contamination from confidence. After illustrating the underlying principles and mechanisms of LogProber, we first validate it through dedicated experiments in which we fine-tune a LLM (LLaMA; Touvron et al. (2023)) on a small set of cognitive question items taken from a recent study Yax et al. (2024a).
The results show that our method is effective in detecting contamination when it involves both the question and its answer. However, in additional model training experiments, we also highlight the inherent limitations of the algorithm in cases where contamination involves only the answer. Finally, we extend our analysis with another set of dedicated experiments using a wide range of questions from standard benchmarks (MMLU Hendrycks et al. (2021) and BigCodeBench Zhuo et al. (2025)), compare LogProber’s performance to that of CDD, and conclude by discussing their relative advantages and weaknesses across different types of contamination.
@@ 2.1 What is contamination and how does it relate to confidence?
Benchmarks typically consist of collections of questions (’Q’) paired with correct answers (’A’). LLM performance is usually assessed by providing ’Q’ as input (e.g., as context or a prompt) and evaluating the model’s generated response. This response is then compared to the known correct answer (’A’) to determine the model’s accuracy.
Contamination broadly refers to the presence of testing materials (benchmark questions or cognitive task items) in the training corpus. In such cases, the LLM’s performance may be attributed directly to memorization of the training data, rather than to genuine generalization or emergent cognitive capabilities. In the context of benchmarks and cognitive tasks, the key question becomes whether the model was trained on the test material—specifically, on the ’Q-A’ pairs. If a full ’Q-A’ sample appears in the training data, the model may learn to predict ’A’ from ’Q’ automatically, without demonstrating real "understanding" (or rather generalization).
In the first part of this paper, we assume that the model has been trained on the complete ’Q-A’ sequence, as opposed to having encountered only the question ’Q’ or the answer ’A’ in isolation. Later in the paper, we will challenge this assumption, exploring other forms of contamination.
Contamination (i.e., the presence of a given ’Q-A’ pair in the training corpus) makes the generation of the response ’A’ following the context ’Q’ highly probable (this is the logic behind the CDD method). However, contamination is not the only reason a LLM might produce a highly probable response. In fact, by analogy with human cognitive science, a LLM’s response probability to a given query is generally interpreted as a reflection of the model’s confidence Efklides (2006). It is therefore important to clarify the distinction and relationship between the key concepts of contamination and confidence. Contamination refers to a property of the model — said to be contaminated — arising from a leakage of the testing material in its training corpus. Confidence refers to a property of the model’s response to a given input — namely, that the model’s answer is confident. Crucially, while responses to contaminated ’Q-A’ pairs will generally appear highly confident, high confidence can also be achieved in other ways. To illustrate this, consider prompting a LLM with the following question:
"What is the next character in the sequence 17817817817817817?"
Many models will complete the sequence with "8" with very high probability (i.e., confidence), due to their well-demonstrated inferential and in-context learning abilities Brown et al. (2020a). However, it is highly unlikely that this specific ’Q-A’ pair was present in the training corpus.
@@ 2.2 A way to measure contamination in LLMs: the LogProber algorithm
Going back to our goal of finding contamination in LLMs, most models are accessed in a ’black box’ setting, where we do not have access to the training corpus nor the model’s weights; we can only infer contamination from the probabilities of the generated tokens. If the sequence ’Q-A’ appears in the training data, the model is contaminated by it, thus the probability of responding ’A’ from ’Q’ should be high. Such probability could be estimated either by directly accessing the tokens’ logprobabilities or by running multiple queries and deriving "empirical" distribution of ’A’. Then, as proposed in Dong et al. (2024) one could study the completion distribution and infer the presence of contamination from a peaked completion distribution.
@@
However, this approach does not rule out that a LLM responds ’A’ to a given ’Q’ with high confidence because of its inferential and generalization capabilities. Indeed the ’A’ token probabilities mix confidence and contamination behaviours making the measure unreliable to disentangle between the two. To tackle this issue, our approach proposes an alternative method that switches the focus on ’Q’ instead of ’A’. Indeed, while it is possible to find a peak in the generation distribution of ’A’ due to a high confidence (and not because of contamination), it is more unlikely that the model is able to predict the question ’Q’ confidently without having been explicitly trained on it.
@@
Assuming we can access the probability of each consecutive token in a given ’Q’ sequence, we can plot the cumulative log-probability of this sequence. This graph can be understood as the ’surprise’ the model has when generating a new token after all the previous ones. Crucially, for ’Q’ present in the corpus we can expect the probability of the subsequent tokens to approach p=1 as soon as the LLM recognizes the sentence and therefore the cumulative log-probability will quickly achieve a plateau. To understand this point see for example what the shape of a very frequent phrase (e.g., the first line of a very famous and old poem Divine Comedy " When half way through the journey of our life"; Figure 1) should look like compared to a (semantically equivalent), but not as frequent, sentence ("Once at the middle of the voyage of my life" Figure 1; right). In the second case, successive tokens will be comparatively more surprising (see 1).
Once the log-probability curves of ’Q’ sequences are obtained, we can analyze their shape in order to quantify how much they plateau to get insight into whether ’Q’ was present in the corpus used to train the model. To measure this plateau in the cumulative logprobability graph, we propose to sort these cumulative logprobability values in increasing order. In this way, if the graph plateaus a lot, the first part of this list should have very low values, and only the very end should have high values, leading to a very small area under the curve. On the other hand, if the graph does not plateau and keeps increasing steadily, the area under the curve should be much larger. As such, the algorithm we propose here consists of sorting the cumulative logprobabilities in increasing order and estimating the area under the curve by summing the sorted values (step function approximation of the integral). In practice, this integral should have a lot of variance between contaminated and uncontaminated sequences. Thus, we will work with the logarithm of this integral instead that we call Safe Score. Experimentally we found that if the Safe Score is lower than 1, the item is likely contaminated. The algorithm is further detailed in Appendix C and is illustrated in Figure 1.
@@ 2.3 Predictions and training experiment
To test the validity of this method, we designed several experiments. We first show the algorithm works to spot contamination in LLMs and we then compare it with CDD Dong et al. (2024) on tasks where the model is susceptible to exhibit high confidence and then on more general topics.
In the first experiment, we compared the Safe Score of LogProber obtained by fitting ’Q’ in a cognitive psychology test. Despite not being well known in the machine learning literature, these tests are particularly interesting when testing contamination as they are rather short, there are very few of them but each of them has been used very much in the literature and on forums so that most LLMs know them by heart. We decided to use one of the most famous : the Cognitive Reflection Test (noted as oldCRT; Toplak et al., 2013; Frederick, 2005). To compare the results obtained on these very likely known items, we used a version designed for LLMs published recently with reframed items that models are not familiar with (noted as newCRT; Yax et al., 2024a) (see Table 2 in appendix B). To be sure the LLM used is not familiar with these new items, we decided to focus on Llama 1 7B Touvron et al. (2023) that was trained well before the publication of these reframed items. Secondly, after testing LogProber in the wild, we test it in a controlled environment by contaminating a LLM on purpose on the newCRT items. Thirdly, we explore more fine-tuning scenarios by testing situations where the model is not trained on the ’Q-A’ sequence but only on the question (noted ’Q’ scenario) or only to produce the correct answer to a given question (noted ’A’ scenario). These experiments further allowed us to explore the complex relationship between contamination and the performance of a model. Finally, we compare LogProber with CDD Dong et al. (2024) first in a MCQ task with a top-end language model likely to be very confident on this task. Then we compare both in a less specific scenario.
@@
| STD-Training | is the original Alpaca training Taori et al. (2023) with no data from new CRT. We only simplified the instruction prompting (see Appendix A). | Q: A bat and a ball cost £1.10 in total. The bat costs £1.00 more than the ball. How much does the ball cost? A: The ball costs £0.05. |
| QA-Training | involves the alpaca training with new CRT pairs of questions and answers. Thus the model is trained to predict each token of the question as well as each token of the answer given the question. | Q: A bat and a ball cost £1.10 in total. The bat costs £1.00 more than the ball. How much does the ball cost? A: The ball costs £0.05. |
@@
| Q-Training | only learns the question’s tokens and therefore is only trained to predict the question’s tokens and not the answer. | Q: A bat and a ball cost £1.10 in total. The bat costs £1.00 more than the ball. How much does the ball cost? A: The ball costs £0.05. |